# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ Appendix D On previous proposals of the Sheaf Hypergraph Laplacian
Let $\mathcal{H}=(V,E)$ be a hypergraph with hyperedge degrees $\delta_{e}$ and let $\mathcal{F}_{v\trianglelefteq e}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be linear restriction maps from node $v$ to hyperedge $e$. The Laplacian $\mathcal{L}_{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ has $d\times d$ blocks indexed by $u,v\in V$:
$(\mathcal{L}_{\mathcal{F}})_{uu}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\qquad(\mathcal{L}_{\mathcal{F}})_{uv}=-\!\!\!\sum_{\begin{subarray}{c}\\ e:u,v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$
@@
$\bigl(\mathcal{L}_{\mathcal{F}}(x)\bigr)_{u}$ $=\sum_{v\in V}(\mathcal{L}_{\mathcal{F}})_{uv}\,x_{v}$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e}\,x_{u}-\sum_{e:\,u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\,x_{v}$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\left(\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e}\,x_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\,x_{v}\right)$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\left(\mathcal{F}_{u\trianglelefteq e}\,x_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\mathcal{F}_{v\trianglelefteq e}\,x_{v}\right).$
@@
$\bigl(\mathcal{L}_{\mathcal{F}}(x)\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\mathcal{F}_{u\trianglelefteq e}\,x_{u}-\mathcal{F}_{v\trianglelefteq e}\,x_{v}\right).$
@@
$(\mathcal{L}_{\mathcal{F}})_{v_{1}v_{1}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{1}\trianglelefteq e_{1}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{2}v_{2}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{1}}+\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{2}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{3}v_{3}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}}+\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{2}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{4}v_{4}}$ $=\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{4}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{1}v_{2}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{1}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{1}v_{3}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{1}v_{4}}$ $=0,$
$(\mathcal{L}_{\mathcal{F}})_{v_{2}v_{3}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}}-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{2}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{2}v_{4}}$ $=-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}},$
$(\mathcal{L}_{\mathcal{F}})_{v_{3}v_{4}}$ $=-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}}.$
@@
$\operatorname{eig}\left(\mathcal{L_{\mathcal{F}}}^{\mathrm{}}\right)=\left\{-2,\ \tfrac{2}{3},\ \tfrac{4}{3},\ 2\right\}.$
@@
Since a negative eigenvalue appears, $\mathcal{L}_{\mathcal{F}}$ is not positive semidefinite in this example.
Appendix E Implementation details
E.1 Computational Complexity
Comparison between DSHN and DSHNLight
Table 5: DSHN vs DSHNLight- FLOPS and Step Time (in ms) Analysis Across Different Datasets (Mean $\pm$ Standard Deviation)
Avg FLOPs/epoch()($\downarrow$) Avg step time $(\downarrow)$
Dataset DSHN DSHNLight DSHN DSHNLight
Roman OOM 12,898,147,996,391 $\pm$ 43,606 OOM 1050.30 $\pm$ 152.04
Amazon OOM 15,061,770,374,298 $\pm$ 0 OOM 1080.26 $\pm$ 159.91
Comparison between DSHN and other models
Fig. 6 reports the average test accuracy of five representative models under approximately the same parameter budget. The results indicate that model size alone does not explain the performance of DSHN. For instance, although SheafHyperGNN and EDHNN have a comparable number of parameters, their accuracy is significantly lower, being these undirected methods. In contrast, DSHN achieves an improvement of about 8% over the strongest directed baselines, despite having the same number of parameters thanks to the expressive power associated to complex-valued and directional restriction maps.
Asymptotic Complexity
  1. 1.
    Linear Transformation. The feature transformation is defined as
    $X^{\prime}=(I\otimes W_{1})XW_{2}$
    where $W_{1}\in\mathbb{R}^{d\times d}$ and $W_{2}\in\mathbb{R}^{f\times f}$. The resulting complexity is $\mathcal{O}(n(d^{2}f+df^{2}))=\mathcal{O}(n(cd+cf))=\mathcal{O}(nc^{2})$, where $c=df$.
  2. 2.
    $Q^{\vec{\mathcal{F}}}_{N}X^{\prime}.$
    The sparsity pattern of $Q^{\vec{\mathcal{F}}}_{N}$ comes directly from the incidence matrix: each hyperedge of size $|e|$ induces $|e|^{2}$ nonzero blocks through the outer product $B^{(q)}(e,:)^{\dagger}B^{(q)}(e,:)$. Summing across all hyperedges gives a total of $\mathsf{S}_{2}=\sum_{e\in\mathcal{E}}|e|^{2}=\mathcal{O}\big(m\bar{e}^{2})$ nonzero blocks, where $\bar{e}$ is the average hyperedge size [^2] One could also upper bound the $\mathsf{S}_{2}$ term with $\mathcal{O}\big(m{n}^{2})$, however, that approximation would be highly pessimistic, considering a fully-dense representation of the hypergraph, where each hyperedge connects all nodes.. Applying the Laplacian then requires $\mathcal{O}(m\bar{e}^{2}c)$ for diagonal maps and $\mathcal{O}(m\bar{e}^{2}dc)$ with non-diagonal maps,
  3. 3.
    Learning the Sheaf. Restriction maps are predicted as
    $\Phi(x_{v},x_{e})=\sigma\!\left(V\bigl(x_{v}\,\|\,x_{e}\bigr)\right)$
    where $V$ is a learnable transformation and $\sigma$ a nonlinearity. The resulting $f$-dimensional vector is then used as input to $V$ for every node-edge incidence. Indicating with $\bar{v}$ the average number of partecipations of a node to an hyperedge, the computational complexity is $\mathcal{O}(\bar{v}mc)$ in the diagonal case, and $\mathcal{O}(\bar{v}md^{2}c)$ in the non-diagonal case.
  4. 4.
    Constructing the Laplacian. In the hypergraph setting we assemble
    $Q^{\vec{\mathcal{F}}}_{N}=D_{V}^{-\frac{1}{2}}{B^{(q)}}^{\dagger}D_{E}^{-1}B^{(q)}D_{V}^{-\frac{1}{2}}.$
    The work naturally splits into two steps:
    1. (a)
      Degree normalization. This involves computing the node and hyperedge degree matrices, $D_{V}^{-\tfrac{1}{2}}$ and $D_{E}^{-1}$. For vertex degree normalization each node requires aggregating contributions from its incident hyperedges, giving $\mathcal{O}(m\bar{e}d)$ operations in the diagonal case and $\mathcal{O}(m\bar{e}d^{3})$ in the non-diagonal case (since each block is $d\times d$), to which it must be added the cost of inverting the block-diagonal matrices, adding to the complexity $\mathcal{O}\big(nd)$ in the diagonal case and $\mathcal{O}\big(nd^{3})$ in the non-diagonal case while since $D_{E}^{-1}$ is obtained by expanding to matrix for the scalar hyperedge degrees $\delta_{e}$ this cost adds a trascurable term to the asymptotic complexity.
    2. (b)
      Sparse product. Forming the term
      $Q^{\vec{\mathcal{F}}}_{N}=D_{V}^{-\frac{1}{2}}{B^{(q)}}^{\dagger}D_{E}^{-1}B^{(q)}D_{V}^{-\frac{1}{2}}.$
      requires, for each hyperedge $e$, generating block interactions among all pairs of nodes it contains. This gives a total of $\mathsf{S}_{2}=\sum_{e\in\mathcal{E}}|e|^{2}$ block products. The cost is $\mathcal{O}(\mathsf{S}_{2}d)$ in the diagonal case and $\mathcal{O}(\mathsf{S}_{2}d^{3})$ in the non-diagonal case. Since the normalization terms $D_{V}^{-\tfrac{1}{2}}$ are block-diagonal operations they do not contribute substantially in the overall complexity. Since $\mathsf{S}_{2}=\mathcal{O}\big(m\bar{e}^{2}\big)$, the dominant cost becomes $\mathcal{O}(m\bar{e}^{2}d)$ for diagonal maps and $\mathcal{O}(m\bar{e}^{2}d^{3})$ for non-diagonal maps.
By summing the overall contributions we get: $\quad\mathcal{O}\!\left(n\,(c^{2}+d)\;+\;m\,(\,\bar{e}d+\bar{e}^{2}(d+c)+\bar{v}\,c\,)\right)$ in the diagonal case and $\quad\mathcal{O}\!\left(n\,(c^{2}+d^{3})+m\,(\,\bar{e}d^{3}+\bar{e}^{2}(d^{3}+dc)+\bar{v}\,d^{2}c\,)\right)$ in the non-diagonal case.
Considerations on the Asymptotic Complexity
E.2 Architectural Choices
Layer Normalization
We apply layer normalization to the learnable parameters $W_{1}\in\mathbb{R}^{d\times d}$ and $W_{2}\in\mathbb{R}^{f\times f}$. Since the input signal to each convolutional layer is complex-valued, we adopt a complex normalization strategy as proposed in Trabelsi et al. [2018], Barrachina et al. [2023], where each complex feature is treated as a two-dimensional real vector $(\Re(x),\Im(x))$. Specifically, we compute the full $2\times 2$ covariance matrix:
$x_{o}=\gamma\tilde{x}+\beta,$
Residual Connections
Following observations from Bodnar et al. [2022], we include residual connections in our convolutional layers, which we found to help the architecture in certain datasets. The use of residuals is treated as a tunable hyperparameter (see Section C.3). With this addition, a convolutional layer takes the form:
$X_{t+1}=\sigma\left(Q^{\vec{\mathcal{F}}}_{N}\,(I_{n}\otimes W_{1})\,X_{t}\,W_{2}+X_{t}\right)\in\mathbb{C}^{nd\times f}.$
Activation Function
For the activation function, we adopt the complex ReLU commonly employed in related works [Zhang et al., 2021, Fiorini et al., 2023, 2024]. It is defined as:
DSHNLight

### v2

@@ Appendix E On previous proposals of the Sheaf Hypergraph Laplacian
Let $\mathcal{H}=(V,E)$ be a hypergraph with hyperedge degrees $\delta_{e}$ and let $\mathcal{F}_{v\trianglelefteq e}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be linear restriction maps from node $v$ to hyperedge $e$. The Laplacian $\mathbf{L}^{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ has $d\times d$ blocks indexed by $u,v\in V$:
$(\mathbf{L}^{\mathcal{F}})_{uu}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e},\qquad(\mathbf{L}^{\mathcal{F}})_{uv}=-\!\!\!\sum_{\begin{subarray}{c}\\ e:u,v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}$
@@
$\bigl(\mathbf{L}^{\mathcal{F}}(\mathbf{x})\bigr)_{u}$ $=\sum_{v\in V}(\mathbf{L}^{\mathcal{F}})_{uv}\,\mathbf{x}_{v}$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{e:\,u\in e}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\,\mathbf{x}_{v}$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\left(\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\mathcal{F}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right)$
$=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\left(\mathcal{F}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\mathcal{F}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$
@@
$\bigl(\mathbf{L}^{\mathcal{F}}(\mathbf{x})\bigr)_{u}=\sum_{e:\,u\in e}\frac{1}{\delta_{e}}\,\mathcal{F}_{u\trianglelefteq e}^{\top}\sum_{\begin{subarray}{c}v\in e\\ v\neq u\end{subarray}}\left(\mathcal{F}_{u\trianglelefteq e}\,\mathbf{x}_{u}-\mathcal{F}_{v\trianglelefteq e}\,\mathbf{x}_{v}\right).$
@@
$(\mathbf{L}^{\mathcal{F}})_{v_{1}v_{1}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{1}\trianglelefteq e_{1}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{2}v_{2}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{1}}+\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{2}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{3}v_{3}}$ $=\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}}+\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{2}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{4}v_{4}}$ $=\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{4}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{1}v_{2}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{2}\trianglelefteq e_{1}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{1}v_{3}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{1}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{1}v_{4}}$ $=0,$
$(\mathbf{L}^{\mathcal{F}})_{v_{2}v_{3}}$ $=-\tfrac{1}{\delta_{e_{1}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{1}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{1}}-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{3}\trianglelefteq e_{2}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{2}v_{4}}$ $=-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{2}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}},$
$(\mathbf{L}^{\mathcal{F}})_{v_{3}v_{4}}$ $=-\tfrac{1}{\delta_{e_{2}}}\,\mathcal{F}_{v_{3}\trianglelefteq e_{2}}^{\top}\mathcal{F}_{v_{4}\trianglelefteq e_{2}}.$
@@
$\operatorname{eig}\left(\mathbf{L^{\mathcal{F}}}\right)=\left\{\tfrac{4}{3},\ \tfrac{1}{3},\ \tfrac{1+\sqrt{17}}{6},\ \tfrac{1-\sqrt{17}}{6}\right\}.$
@@
Since a negative eigenvalue appears, $\mathbf{L}^{\mathcal{F}}$ is not positive semidefinite in this example.